# SweetTokenizer: Semantic-Aware Spatial-Temporal Tokenizer for Compact Visual Discretization

### v2

@@ 3.2.3 Language-based Latent Codebook
Previous works [29, 54, 57] have shown that text representations can enhance image VQ-VAEs , as the text provides additional semantic information from pre-trained language models. However, previous works mainly focus on the relationship between static image appearance and text semantics. Our experiment in Table 5 shows this is insufficient for video data, as static and motion information are typically embedded in different subsets of vocabularies.
To address this, we construct two separate codebooks, the spatial quantization codebook and the temporal quantization codebook, according to vocabulary attributes, as illustrated in Figure 2. We first extract candidate vocabularies from video captions, which is obtained by Qwen-2.5B [1] instruct model. We filter out words that occur with low frequency. Afterward, we extract Qwen-2.5B [1] text embedding of these vocabularies to fill in the columns of our codebook $C\in\mathbb{R}^{L\times D}$ and split it into four subsets: nouns, adjectives, verbs, and adverbs. Note that $C_{noun},C_{adj}$ are for spatial latent tokens $\Omega_{s}$ while $C_{verb},C_{adv}$ are for temporal latent tokens $\Omega_{t}$.
@@
Given two encoded continuous latent vectors: $z_{s}\in\Omega_{s}$ and $z_{t}\in\Omega_{t}$, $z_{s}$ is passed through spatial quantization codebook, and $z_{t}$ is passed through temporal quantization codebook. The quantized $\hat{z}_{s}$ and $\hat{z}_{t}$ are obtained by nearest neighbor searching:
@@
$\hat{z}_{s}=\mathcal{P}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{noun}\cup C_{adj}}||z_{s}-\mathcal{P}(c_{i})||,$ (3)
$\hat{z}_{t}=\mathcal{P}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{verb}\cup C_{adv}}||z_{t}-\mathcal{P}(c_{i})||.$ (4)
To maintain text semantic information, we freeze the codebook and use a tiny projector network $\mathcal{P}$ to map from text space into visual space, following the strategy in [57]. Finally, the gradient is passed to the encoder via vector-quantization commitment loss proposed in [41], a common method to approximate differentiability:
@@
where $sg[\cdot]$ is stop-gradient operator. Figure 4 also shows that the encoded latent words by SweetTokenizer capture semantic meanings related to both visual appearance and motion.
@@
3.2.4 Progressive Training Schedule
To ensure better convergence, we follow the philosophy of curriculum learning [2], thus designing a progressive training schedule comprising three stages.
We adopt the proxy codes strategy following [56] to leverage pre-trained visual priors. Assuming $\Psi_{e}(\cdot)$ is a pre-trained visual tokenizer and $\Psi_{d}(\cdot)$ is a pre-trained pixel decoder. $\Psi_{e}(\cdot)$ maps visual inputs $x$ into well-trained integer token indices $\Psi_{e}(x)$, which serves as a proxy code classification label of our CQAE model, similar to the knowledge distilling approach [16]. The motivation is that we could bypass the early oscillatory training phase of GAN loss [13] as demonstrated in [56, 11], which allows us to apply it later once the training becomes stable. In practice, we use OmniTokenizer [45]’s encoder and decoder as $\Psi_{e}(\cdot)$ and $\Psi_{d}(\cdot)$.
First, we train $CQAE_{s}$ on spatial patches $v_{s}$ using image data only as pre-training in Stage 1, with the supervision of proxy code:
$\mathcal{L}_{stage1}=\mathcal{L}_{c}(CQAE_{s}(v_{s}),\Psi_{e}(x))+\mathcal{L}_{vq},$ (6)
where $\mathcal{L}_{c}$ denotes cross entropy loss and $\mathcal{L}_{vq}$ denotes codebook commitment loss in Eq 5. Note that SweetTokenizer adopts spatial-temporal disentanglement, so the model weights of $CQAE_{s}$ can be later adapted to image-level tasks via finetuning. Experiments in Table 3 show that our spatial branch already achieves satisfactory image reconstruction results.
Second, we jointly train $CQAE_{s},CQAE_{t}$ on $v_{s},v_{t}$ using video data in Stage 2. Since $CQAE_{s}$ is well-initialized, $CQAE_{t}$ can focus on learning temporal motions:
$\mathcal{L}_{stage2}=\mathcal{L}_{c}(CQAE_{s}(v_{s}),\Psi_{e}(x))+$ (7)
$\quad\quad\quad\quad\mathcal{L}_{c}(CQAE_{t}(v_{t}),\Psi_{e}(x))+\mathcal{L}_{vq}.$
Finally, in Stage 3, we combine the pre-trained pixel decoder $\Psi_{d}(\cdot)$ together with our CQAE and finetune them by pixel-level losses in Stage 3 to boost final performance:
$\mathcal{L}_{stage3}=\mathcal{L}_{(m,g,p)}(\Psi_{d}(CQAE(v)),x),$ (8)
where $\mathcal{L}_{m}$ denotes mean-squared error loss, $\mathcal{L}_{g}$ denotes GAN loss [13] and $\mathcal{L}_{p}$ denotes perceptual loss [20].
@@ 4.1 Experiments Settings
Tokenizer #Tokens rFVD $\downarrow$
UCF-101 K-600
MaskGIT [6] 4352 240 202
VQGAN [11] 4352 299 270
TATS [12] 4096 162 -
MAGVIT [53] 4096 58 -
OmniTokenizer [45] 5120 42 26
SweetTokenizer 5120 18 8
SweetTokenizer 256 + 1024 44 28
Table 1: Video reconstruction FVD on the UCF-101 and K-600 dataset, using a frame resolution 256 × 256. $\ast$” denotes training SweetTokenizer without token compression.
Tokenizer #Params #Tokens gFVD $\downarrow$
CogVideo [17] 9.4B 6800 626
TATS [12] 321M 4096 332
Video-LaVIT [18] 7B 512 280
OmniTokenizer [45] 650M 5120 191
SweetTokenizer 650M 5120 154
SweetTokenizer 650M 256 + 1024 128
Table 2: Comparison of class-conditional generation results on UCF-101. Each video is composed of 17 frames with a resolution of 256 $\times$ 256. We only compare autoregressive generators (AR) as baselines. “#Params” refers the parameter number of generator.
@@
We evaluate the tokenization performance of SweetTokenizer on image and video datasets, including ImageNet [8], UCF-101 [34], and Kinetics-600 [21, 5]. Following [45], all images and video frames are resized to 256 × 256 resolution for experiments. The semantic capabilities of SweetTokenizer are tested through few-shot image classification on Real-Name Open-Ended miniImageNet [39] and few-shot video action recognition on UCF-101, as described in [58].
@@
SweetTokenizer 1024 10481 0.37
Table 3: Image reconstruction FID on the ImageNet dataset, using a resolution of 256 $\times$ 256. $\diamond$” denotes continuous latent space without quantization. $\ast$” denotes training SweetTokenizer without token compression.
Implementation Details.
SweetTokenizer adopts a spatial-temporal architecture consisting of 8 spatial layers and 4 temporal layers, with both the encoder and decoder configured to a hidden dimension of 512. The latent space dimension is set to 256. For the LLM codebook quantizer, we exclude words with a frequency below 5, resulting in a selection of 5,078 nouns, 5,403 adjectives, 9,267 verbs, and 1,872 adverbs. This forms a spatial codebook of size 10,481 and a temporal codebook of size 11,139. We utilize a graph convolution network (GCN) to project Qwen-2.5 embeddings [1] into the visual latent space. Graph edges are constructed when a pair of words co-occur within a 5-token window in the text. The model is trained with a progressive three-stage schedule: Stage 1 uses a batch size of 32 for 300K iterations, Stage 2 uses a batch size of 8 for 300K iterations, and Stage 3 uses a batch size of 4 for 500K iterations. All training is performed on NVIDIA A100 GPUs. Adam [23] is employed for optimization ($\beta_{1}=0.9$ and $\beta_{2}=0.99$). During each stage, we use a cosine learning rate scheduler with a max learning rate of 1e-4 and a min learning rate of 1e-5, warmed up by 10K iterations.
OmniTokenizer [45]
SweetTokenizer (ours)
Figure 5: Class-conditional video generation result using OmniTokenizer [45] and our SweetTokenizer, on UCF-101 dataset. The action class label of each row is: “BlowingCandles”, “PlayingTabla”, “TennisSwing”, “PlayingViolin” and “HorseRiding”.
@@ 4.2 Video Reconstruction & Generation
We first evaluate the tokenization capability of SweetTokenizer on the UCF-101 and K-600 video datasets. As shown in Table 1, SweetTokenizer uses only 1,280 tokens (256 spatial tokens and 1,024 temporal tokens), which is four times fewer than OmniTokenizer’s 5,120 tokens. Despite this significant reduction, it achieves comparable performance, with rFVD scores of 44.35 on UCF-101 and 27.95 on K-600, demonstrating its effectiveness in compressing tokens while maintaining high fidelity. Notably, at the same compression ratio, SweetTokenizer significantly outperforms all baselines, achieving an rFVD of 18.74 on UCF-101, 57.1% lower than OmniTokenizer’s 42.35, and an rFVD of 7.51 on K-600, 71.0% lower than OmniTokenizer’s 25.97.
The generative capability of SweetTokenizer is evaluated on UCF-101 in a class-conditional generation task. SweetTokenizer is used to extract quantized spatial tokens ($\hat{\Omega}_{s}$) and temporal tokens ($\hat{\Omega}_{t}$) from UCF videos. These tokens are then concatenated to form training sequences for a 24-layer autoregressive (AR) transformer [49], following the same generation protocol as OmniTokenizer. As shown in Table 2, SweetTokenizer achieves a significant performance improvement, with a gFVD score of 128, 32.9% lower than OmniTokenizer’s 191 gFVD. This improvement is attributed to SweetTokenizer’s effective token compression, which substantially reduces the training complexity for downstream autoregressive models.
The visualization results are presented in Figure 5. To ensure a fair comparison, we select generated videos with similar appearances, as the generation process inherently involves randomness. The results demonstrate significantly improved detail, such as clearer human facial features and finer table textures. Additionally, SweetTokenizer effectively preserves temporal consistency, even under large motion scenarios. Further quantitative and qualitative results on video datasets are provided in the supplementary.
@@ 4.3 Image Reconstruction
We evaluate the image tokenization performance of SweetTokenizer on ImageNet by fine-tuning the spatial branch $CQAE_{s}$ with a pixel-level loss during Stage 3. As shown in Table 3, we compare SweetTokenizer with recent methods under various token compression settings. With 256 spatial tokens, SweetTokenizer outperforms TiTok [56] by 27.8%, reducing rFID from 1.01 to 0.73. When using 1,024 spatial tokens, SweetTokenizer achieves a significant improvement over both VQ-based and non-VQ-based methods (marked $\diamond$), achieving an rFID of 0.37, which surpasses LlammaGen-8 [35] by 37.3%. More quantitative and qualitative results on image reconstruction and generation tasks are shown in the supplementary.
@@ 4.4 Ablation Studies
Spatial-Temporal Decoupling.
We demonstrate that naively flattening video tokens into a 1D sequence is infeasible, as shown in Table 4. Training SweetTokenizer without decoupling results in significantly degraded performance. We attribute this degradation to two factors: (1) the flattening operation discards substantial consecutive temporal information, and (2) without decoupling, the model cannot leverage the crucial spatial pretraining from large-scale image datasets, as further discussed in the next paragraph.
Progressive Training Schedule.
This paragraph highlights the advantages of our progressive training schedule for the model, evaluated on the UCF-101 dataset. The results, presented in Table 4, demonstrate the critical role of each training stage. First, we underscore the importance of spatial training: omitting Stage 1 results in significantly poor rFVD scores of 269.34 and 501.12 for the combinations of “Stage 2 + Stage 3” and “Stage 2” alone, respectively. Secondly, the proxy code finetuning strategy employed in Stage 3 has a substantial impact on overall performance, improving the rFVD score from 155.92 with “Stage 1 + Stage 2” to 44.35 with the full sequence of “Stage 1 + Stage 2 + Stage 3”. These findings highlight the necessity of our complete progressive learning schedule, as each stage contributes to enhancing the model performance.
Architecture & Training Schedule rFVD $\downarrow$
Stage2 w/o decoupling 892.70
Stage2 501.12
Stage2 $+$ Stage3 269.34
Stage1 $+$ Stage2 155.92
Stage1 $+$ Stage2 $+$ Stage3 (SweetTokenizer) 44.35
Table 4: Ablation study of spatial-temporal decoupling and progessive training schedule.
UCF-101 rFVD $\downarrow$
Baseline (w/o LLC) 67.94
$+$ Spatial LLC 55.11
$+$ Spatial & Temporal LLC (SweetTokenizer) 44.35
$+$ CLIP [31]-based LLC 45.56
$+$ Qwen [1]-based LLC (SweetTokenizer) 44.35
ImageNet rFID $\downarrow$
Baseline (w/o LLC) 1.26
$+$ Spatial LLC (SweetTokenizer) 0.73
Table 5: Ablation study of language-based latent codebook (LLC) settings on UCF-101 and ImageNet.
Language-based Latent Codebook.
We evaluate the impact of our architecture design on model performance, focusing on the effects of the spatial-temporal language-based latent codebooks (LLC) on the ImageNet and UCF-101 datasets. As illustrated in Table 5, the integration of the spatial language codebook consistently enhances performance across both video and image modalities, improving the rFVD score from 67.94 to 55.11 and the rFID score from 1.26 to 0.73. Moreover, our temporal language codebook significantly benefits video reconstruction tasks, further reducing the rFVD score from 55.11 to 44.35. This underscores the importance of our unique design for video modalities. We attribute these improvements to the robust semantic features provided by language-based embeddings and their inherent alignment between visual and text spaces. Additionally, we compare different types of language-based embeddings, such as using CLIP [31] embeddings in place of Qwen-2.5B [1] embeddings. The experiments indicate that while CLIP embeddings yield satisfactory results, employing a more powerful pre-trained language model leads to superior performance.

### v3

@@ 3.2.3 Motion-enhanced Language Codebook (MLC)
To mitigate information loss during compression, we introduce an language codebook (LC) quantizer to enhance semantic richness. The Previous works [24, 49] have shown that text representations can enhance image VQ-VAEs , as the text provides additional semantic information from pre-trained language models. However, previous works mainly focus on the relationship between static image appearance and text semantics [52]. Our experiment in Table 5 shows is insufficient for video data.
To address this, we propose a Motion-enhanced Language Codebook (MLC), where the video motion information is enhanced via action-related vocabularies. Specifically, we split the dictionary into four subsets: nouns, adjectives, verbs, and adverbs. Intuitively, static and appearance information is typically embedded in nouns and adjectives, while motion information is generally embedded in verbs and adverbs. Therefore, we choose nouns and adjectives for spatial query tokens $\mathbf{Q_{s}}$, and verbs and adverbs for temporal query tokens $\mathbf{Q_{t}}$. Figure 3 also shows that the encoded latent words by SweetTok capture semantic meanings related to both visual appearance and motion.
As for details, we first extract candidate vocabularies of the whole dataset from video captions. Afterward, we extract CLIP [27] text embedding of these vocabularies to fill in the columns of our codebook $C\in\mathbb{R}^{L\times D}$. We utilize a graph convolution network $\mathcal{F}$ to project CLIP embeddings [27] into the visual latent space. Graph edges are constructed when a pair of “spatial-spatial”, “spatial-temporal” or “temporal-temporal” words co-occur within a 5-token window in the current video caption.
@@
Given two encoded continuous latent vectors: $z_{s}\in\mathbf{Q_{s}}$ and $z_{t}\in\mathbf{Q_{t}}$, $z_{s}$ is passed through spatial quantization codebook, and $z_{t}$ is passed through temporal quantization codebook. The quantized $\hat{z}_{s}$ and $\hat{z}_{t}$ are obtained by nearest neighbor searching:
@@
$\hat{z}_{s}=\mathcal{F}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{noun}\cup C_{adj}}||z_{s}-\mathcal{F}(c_{i})||,$ (10)
$\hat{z}_{t}=\mathcal{F}(c_{i}),i=\mathop{\arg\min}\limits_{c_{i}\in C_{verb}\cup C_{adv}}||z_{t}-\mathcal{F}(c_{i})||.$ (11)
Finally, the gradient is passed to the encoder via vector-quantization commitment loss proposed in [36], a common method to approximate differentiability ($sg[\cdot]$ stands for stop-gradient operator):
@@ 4.1 Experiments Settings
We evaluate the tokenization performance of SweetTok on video datasets, including UCF-101 [30], and Kinetics-600 [16, 4]. Following [41], all video frames are resized to 256 $\times$ 256 resolution for experiments. Note that some of the previous works use a resolution of 128 $\times$ 128, which cannot be directly compared to our work due to the difference in task difficulty. However, we still include these results in the table and highlight them in gray. Moreover, we fine-tune SweetTok’s spatial component on ImageNet [7] to obtain a strong image tokenizer. The semantic capabilities of SweetTok are tested through few-shot image classification on Real-Name Open-Ended miniImageNet [34] and few-shot video action recognition on UCF-101, as described in [53].
@@
OmniTok [41]
LARP [39]
SweetTok (ours)
Figure 4: Video reconstruction result on UCF-101 dataset. We also visualize the reconstruction and GT error map, where brighter areas indicate larger errors.
Tokenizer #Tokens #Params rFVD $\downarrow$
Tokenizer UCF-101 K-600
MAGVIT-V2 [50] 1280 307M 8.6 -
LARP-L [39] 1024 193M 20 13
MaskGIT [5] 4352 227M 240 202
VQGAN [10] 4352 227M 299 270
TATS [11] 4096 32M 162 -
MAGVIT [48] 4096 158M 58 -
OmniTok [41] 5120 82.2M 42 26
LARP-B [39] 1024 143M 64 35
LARP-L [39] 1024 193M 35 23
SweetTok 5120 128M 11 8
SweetTok 1280 128M 20 25
Table 1: Video reconstruction FVD on the UCF-101 and K-600 dataset, using a frame resolution 256 × 256. $\ast$” denotes training SweetTok without token compression. Lines in “gray” indicate results at a resolution of 128 $\times$ 128.
Implementation Details.
SweetTok adopts a spatial-temporal architecture consisting of 8 spatial layers and 4 temporal layers, with both the encoder and decoder configured to a hidden dimension of 512. The latent space dimension is set to 256. For the LLM codebook quantizer, we exclude words with a frequency below 5, resulting in a selection of 5,078 nouns, 5,403 adjectives, 9,267 verbs, and 1,872 adverbs. This forms a spatial codebook of size 10,481 and a temporal codebook of size 11,139. The model is trained with a batch size of 8 for 1000K iterations. All training is performed on NVIDIA A100 GPUs. Adam [18] is employed for optimization ($\beta_{1}=0.9$ and $\beta_{2}=0.99$). During each stage, we use a cosine learning rate scheduler with a max learning rate of 1e-4 and a min learning rate of 1e-5, warmed up by 10K iterations.
Tokenizer Type #Tokens #Params gFVD $\downarrow$
Generator
MAGVIT [48] AR 1024 306M 265
MAGVIT-V2 [50] AR 1280 307M 109
MAGVIT [48] MLM 1024 306M 76
MAGVIT-V2 [50] MLM 1280 307M 58
LARP-L [39] AR 1024 632M 57
CogVideo [13] AR 6800 9.4B 626
TATS [11] AR 4096 321M 332
Video-LaVIT [14] AR 512 7B 280
OmniTok [41] AR 5120 650M 191
LARP-L [39] AR 1024 632M 99
SweetTok AR 1280 650M 84
SweetTok AR 1280 1.9B 65
Table 2: Class-conditional video generation results on UCF-101. Each video is composed of 17 frames with a resolution of 256 $\times$ 256. “AR” and “MLM” represents autoregressive and masked-language-modeling generator. Lines in “gray” indicate results at a resolution of 128 $\times$ 128.
@@
SweetTok 1024 10481 0.37
Table 3: Image reconstruction FID on the ImageNet dataset, using a resolution of 256 $\times$ 256. $\diamond$” denotes continuous latent space without quantization. $\ast$” denotes training SweetTok without token compression.
OmniTok [41]
LARP [39]
SweetTok (ours)
Figure 5: Class-conditional video generation result on UCF-101 dataset. The action class label of each row is: “PlayingTabla”, “TennisSwing”, and “HorseRiding”.
@@ 4.2 Video Reconstruction & Generation
We first evaluate the tokenization capability of SweetTok on the UCF-101 and K-600 video datasets. As shown in Table 1, SweetTok uses 1,280 tokens (256 spatial tokens and 1,024 temporal tokens), which is four times fewer than OmniTok’s 5,120 tokens and comparable with LARP’s 1024 tokens. Despite its high compression ratio, our method achieves a 52.3% improvement on UCF-101 and competitive performance on K-600 compared with OmniTok. With a similar token count, SweetTok, despite its smaller model size compared to LARP-L, delivers a 42.8% improvement on UCF-101 and comparable results on K-600. Our SweetTok demonstrates significant performance gains, achieving 68.8% and 28.5% improvements in rFVD on UCF-101 and K-600 datasets, respectively, compared to LARP-B with similar model size. Notably, if we increase the token number to 5,120, SweetTok significantly outperforms all baselines, achieving an rFVD of 10.74 on UCF-101 and 7.51 on K-600.
The generative capability of SweetTok is evaluated on UCF-101 in a class-conditional generation task. SweetTok is used to extract decoupled tokens from UCF videos. These tokens are then concatenated to form training sequences for VideoGPT [44], following the same generation protocol as OmniTok. As shown in Table 2, SweetTok achieves a significant performance improvement, with a gFVD score of 84, 56% lower than OmniTok’s 191 gFVD. This improvement is attributed to SweetTok’s effective token compression, which substantially reduces the training complexity for downstream autoregressive models. With equivalent token counts, SweetTok demonstrates superior performance, achieving a 15.1% improvement in gFVD (84 vs. LARP’s 99) at comparable generator sizes. Furthermore, it exhibits scaling law characteristics, with gFVD improving from 84 to 65 as the model scales up to 1.9B parameters.
The visualization results are presented in Figure 5. To ensure a fair comparison, we select generated videos with similar appearances, as the generation process inherently involves randomness. The results demonstrate significantly improved detail, such as clearer human facial features and finer table textures. Additionally, SweetTok effectively preserves temporal consistency, even under large motion scenarios.
Compression Method #Tokens rFVD $\downarrow$
Vanilla Downsample 1280 227.65
Vanilla Query-based (LARP [39]) 1024 35.15
Decoupled Query-based (DQAE) 1280 20.46
Table 4: Ablation study of different token count compression method for video tokenizers.
@@ 4.3 Image Reconstruction
To demonstrate the flexibility of our decoupled query design, we show that by directly fine-tuning the spatial branch $DQAE_{s}$, SweetTok also achieves advanced performance for image reconstruction on ImageNet. As shown in Table 3, we compare SweetTok with recent methods under various token compression settings. With 256 spatial tokens, SweetTok outperforms TiTok [51] by 27.8%, reducing rFID from 1.01 to 0.73. When using 1,024 spatial tokens, SweetTok achieves a significant improvement over both VQ-based and non-VQ-based methods (marked $\diamond$), achieving an rFID of 0.37, which surpasses LlammaGen-8 [31] by 37.3%. Visualization results are shown in Figure 6, SweetTok maintains better global appearance and local details.
GT
TiTok [51]
OmniTok [41]
SweetTok
Figure 6: Visualization of image reconstruction results.
@@ 4.4 Ablation Studies
Decoupled Query AutoEncoder.
We demonstrate the effectiveness of our spatial-temporal decoupling design for token compression. As shown in Table 4, naively downsampling video tokens by 1D linear-interpolation from 5,120 patches into 1,280 tokens results in poor performance of a rFVD of 227.65. Training SweetTok without decoupling (similary to vanilla query-based method in LARP [39]) results in sub-optimal result, obtaining a rFVD of 35.15. Our proposed decoupling query strategy (DQAE) achieves best result of 20.46 rFVD. We attribute this improvement to two factors: (1) the flattening operation discards substantial consecutive temporal information, and (2) without decoupling, the model struggles to learn efficiently from the intertwined temporal and spatial information.
Motion-enhanced Language Codebook.
We evaluate the impact of our architecture design on model performance, focusing on the effects of the motion-enhanced language codebooks (MLC) on UCF-101 datasets. As illustrated in Table 5, vanilla language codebook design enhances rFVD performance from 29.45 to 24.80, leading to a sub-optimal result. Notably, our motion-enhanced temporal language codebook significantly benefits video reconstruction tasks, further reducing the rFVD score from 24.80 to 20.46. This underscores the importance of our unique design for video modalities. Additionally, we compare different types of language-based embeddings, such as using Qwen-2.5B [1] embeddings in place of CLIP [27] embeddings. The experiments indicate that naively increasing the complexity of pre-trained language model is not cost-effective for SweetTok.